OpenStudy (anonymous):

I need help finding the first derivitave of this.: y= sinx - 1/18sin3x...[0, 2pi]. But it has to look like this in the end : cosx (3/2 - 2/3cos^2x)

OpenStudy (chihiroasleaf):

what do you about finding derivative of trigonometric function? like, what is the derivative of sin x ?

OpenStudy (agent0smith):

\[ y= sinx - \frac{ 1 }{ 18\sin3x }\]or \[ y= sinx - \frac{ 1 }{ 18 }\sin3x\] Which is it?

OpenStudy (anonymous):

This is what was given : y'= cosx - 1/6 cos3x = cosx - 1/6[cos (2x +x)] = cosx - 1/6 [cos2x cosx - sin2x sinx] and then we are supposed to find out the next three steps for which i have (trouble): = cosx - 1/6[ (cos^2x - sin^2x) cosx - 2sinxcosxsinx] = cosx - 1/6 [cos^3x - sin^2x cosx - 2sin^2x cosx] = cosx - 1/6 [2cosx - cosx - cos^3x - 2cosx - 2cos^3x] and then, she gave us this part: = cosx - 1/6 [ 2cos^3x - cosx - 2cosx + 2cos^3x] = cosx - 1/6 [4cos^3x -3cosx] = cosx - 2/3cos^3x + 1/2cosx = 3/2cosx - 2/3cos^3x = cosx ( 3/2 - 2/3cos^2x) cosx= 0. 3/2 - 2/3cos^2x=0 x= ?

OpenStudy (anonymous):

And its the 2nd oneyou wrote

OpenStudy (agent0smith):

The steps where you have trouble look like double angle formulas and such

OpenStudy (anonymous):

Yeah, they are

OpenStudy (agent0smith):

Angle sum formulas and probably double angle formulas... http://www.mathwarehouse.com/trigonometry/identities/angle-sum-formula.php

OpenStudy (anonymous):

this is your question \[\Large y=\sin(x)-\frac{1}{18}\sin(3x)\] did i write it correct ?

OpenStudy (agent0smith):

^ that's correct sami

OpenStudy (anonymous):

Yes you did

OpenStudy (anonymous):

My teacher said to use formulas to substitute the sines.. But i dont know how to do it in three steps...

OpenStudy (agent0smith):

You don't have to do it all in three steps, do it bit by bit if you need to. = cosx - 1/6 [cos2x cosx - sin2x sinx] .......................................step 1 = cosx - 1/6[ (cos^2x - sin^2x) cosx - 2sinxcosxsinx] .................step 2 first replace cos(2x) with (cos²x − sin²x)

OpenStudy (agent0smith):

then replace sin2x with (2sinxcosx)

OpenStudy (anonymous):

Well i say three steps because she only gave us that much space to do it

OpenStudy (agent0smith):

Oh, well you can do it in three if you have to. Just replace them one at a time. Then simplify.

OpenStudy (anonymous):

Andd, i already have that. Im stuck on step 3..

OpenStudy (anonymous):

ok since \[\Large \sin(3x)=3\sin(x)-4\sin^3(x)\] y=sin(x)-1/18(3sin(x)-4sin^3(x) y'=cos(x)-1/18(3cos(x)-12sin^2(x)cos(x)] y'=cos(x)[1-1/18[3-12sin^2x]] y'=cos(x)[1-1/6+2/3sin^2(x)]

OpenStudy (anonymous):

just replace sin^2x with 1-cos^2(x) now you get your answer .

OpenStudy (anonymous):

Okay, wait, did you see my 2nd response? Because if i could do that, i would, but she wants us to go the other route.. I have step 1 and 2.. But im stuck on step 3 ...

OpenStudy (anonymous):

did you get it or should i write it in latex ?

OpenStudy (anonymous):

its the three step processs \[\Large \sin(3x)=3\sin(x)-4\sin^3(x)\] question becomes \[\Large y=\sin(x)-\frac{1}{18}[3\sin(x)-4\sin^3(x)]\] differentiate \[\Large y'=\cos(x)-\frac{1}{18}[3\cos(x)-12\sin^2(x)\cos(x)]\] \[\Large y'=\cos(x)[1-\frac{1}{6}+\frac{2}{3}\sin^2(x)]\] \[\Large y'=\cos(x)[1-\frac{1}{6}+\frac{2}{3}[1-\cos^2(x)]]\] replace sin^2(x)=1-cos^2(x) \[\Large y'=\cos(x)[\frac{3}{2}-\frac{2}{3}\cos^2(x)]\]

OpenStudy (agent0smith):

= cosx - 1/6 [cos^3x - sin^2x cosx - 2sin^2x cosx] ..................step 2 = cosx - 1/6 [2cosx - cosx - cos^3x - 2cosx - 2cos^3x] ...........step 3 ^that's the step you're stuck on, yes? I think she made a lot of use of the identities 1-sin^ x = cos^2 x and 1-cos^2 x = sin^2 x

OpenStudy (anonymous):

yes yes yess thank youu okay now i have question, does the last step you did - does that simplify into what she ave me? : Cosx - 1/6[2cos^3x - cosx - 2cosx + 2cos^3x]

OpenStudy (agent0smith):

Let's check... copied from your earlier post: = cosx - 1/6 [cos^3x - sin^2x cosx - 2sin^2x cosx].............. step 2 = cosx - 1/6 [2cosx - cosx - cos^3x - 2cosx - 2cos^3x] .............. step 3 and then, she gave us this part: = cosx - 1/6 [ 2cos^3x - cosx - 2cosx + 2cos^3x] I'll start from step 2: = cosx - 1/6 [cos^3x - sin^2x cosx - 2sin^2x cosx] = cosx - 1/6 [cos^3x - 3sin^2x cosx] = cosx - 1/6 [cos^3x - 3(1-cos^2x) cosx] = cosx - 1/6 [cos^3x - 3cosx + 3cos^3x ] = cosx - 1/6 [4cos^3x - 3cosx] which is what she has up there. Idk what step 3 was about though, maybe you forgot some parentheses or something, cos it doesn't look right.

OpenStudy (anonymous):

Okay, well thank you, so much :)